OpenStudy (tayjane1997):

Please help algebra 2 question!?

OpenStudy (tayjane1997):

A biologist is comparing the growth of a population of flies per week to the number of flies a lizard will consume per week. She has devised an equation to solve for which day (x) the lizard would be able to eat the entire population. The equation is 3x = 5x − 1. Explain to the biologist how she can solve this on a graph using a system of equations

OpenStudy (obaybk3900):

Simplifying 5x + -1 = 3x + 1 Reorder the terms: -1 + 5x = 3x + 1 Reorder the terms: -1 + 5x = 1 + 3x Solving -1 + 5x = 1 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. -1 + 5x + -3x = 1 + 3x + -3x Combine like terms: 5x + -3x = 2x -1 + 2x = 1 + 3x + -3x Combine like terms: 3x + -3x = 0 -1 + 2x = 1 + 0 -1 + 2x = 1 Add '1' to each side of the equation. -1 + 1 + 2x = 1 + 1 Combine like terms: -1 + 1 = 0 0 + 2x = 1 + 1 2x = 1 + 1 Combine like terms: 1 + 1 = 2 2x = 2 Divide each side by '2'. x = 1 Simplifying x = 1